# A Certifiable Machine Learning-Based Pipeline to Predict Fatigue Life of Aircraft Structures

### v1

@@ Abstract
GAG Ground Air Ground
@@ 1 Introduction
Damage estimation caused by fatigue [1] is a critical aspect of aeronautical engineering, as failure to predict it accurately can lead to catastrophic structural breakdowns and compromise both flight safety and operational efficiency. Likewise, fatigue damage assessment is essential for implementing predictive maintenance strategies that extend the life of aircraft and reduce unplanned downtime. Standard fatigue prediction methods rely on a convoluted and resource-intensive process that combines service history of aircraft of similar structural design with aerodynamic and Finite Element Methods simulations and cumulative damage modeling to guide expert-based decision making (CS-25.571 in [2]). The complexity of the whole process makes the development of tailored maintenance routines for new aircraft operators particularly expensive and time-consuming, involving multidisciplinary teams and a range of iterative procedures.
@@
Interestingly, the advent of Machine Learning (ML) techniques and the rapid incorporation of these tools within the aeronautical engineering community (encompassing fields from aerodynamics [3, 4, 5], stress prediction [6] or aeroacoustics [7, 8] to airflow optimization [9] –see [10] for a recent review–) provides a realistic opportunity to streamline and optimize the procedures for fatigue-related damage prediction, as these techniques potentially offer near-instantaneous inference of fatigue life, once properly trained. Specifically in fatigue-related problems, ML has been recently used for load spectrum reconstruction [11, 12], fatigue life prediction [13, 14], and crack growth modeling [15, 16]. However, most approaches treat these steps separately and rely on hybrid frameworks that combine ML with traditional tools, lacking end-to-end ML-enabled fatigue life prediction pipelines.
@@
At the same time, there is an urge of incorporating fully certifiable ML-based procedures within the aeronautical industry. As a matter of fact, a key consideration for ML in safety-critical applications is the ability to certify the predictive model. Aeronautical authorities such as EASA have outlined the main challenges and goals of these industries for trustworthy ML adoption [17, 18], and recent research has started to propose practical methodologies for statistical validation of ML pipelines in industrial settings [6], encompassing data preparation, model training and interpretability, and multiscale uncertainty quantification.
@@
This work proposes a fully integrated and certifiable ML-based pipeline capable of predicting fatigue life at various locations of an aircraft wing, based solely on flight parameters. The pipeline mostly uses a suite of different multilayer perceptrons (MLPs) to approximate different input-output relationships within the fatigue damage assessment. An MLP is an archetypal method in supervised machine learning [19] used to infer (regress or classify) complex input-output representations $x\to y$, where $x\in\mathbb{R}^{m}$ and $y$ is usually another vector for regression tasks or an element from a discrete set in classification tasks. Often visualized as an (artificial neural) network, an MLP consists in a (deep) feedforward neural network representing an overparameterized nonlinear function $\mathscr{F}(x;{\Omega})$, where $\Omega=\{w_{k}\}$ –the set of trainable parameters– identify with the weights of the links between neurons (see [19] for details). The precise architecture of the MLP (number of neurons per layer, number of layers) as well as the precise optimization scheme to train the network are defined in a case-by-case basis.
@@
The rest of this paper is structured as follows: in Section 2, we present the full methodology: we define the physical problem under study, the dataset, the set of ML techniques and software tools, and the statistical techniques used for a thorough validation. Then, in Sec. 3 we present the results, which we anticipate yield a successful ML-driven fatigue life prediction tool. Finally, conclusions and future directions are discussed in Sec. 4.
@@ 2 Methodology
In this section, we present the methodology adopted in this work. First, in subsection 2.1, we describe the physical problem being modeled using machine learning techniques and explain how it is decomposed into smaller, learnable subproblems. Then, in subsection 2.3, we detail the available dataset and the data-splitting strategy used for training, which plays a crucial role in the performance of the models. Finally, in subsection 2.4, we add a few comments on the ML and statistical software libraries we have used.
@@ 2.1 Breaking down the fatigue damage assessment
The goal of this study is to develop a machine learning pipeline to predict fatigue life (expressed as the number of flights to failure) at multiple wing locations –known as principal structural elements (PSEs)–, when flying a specific mission (i.e., a flight characterized by specific flight parameters and operational objectives). In what follows we unfold the process and breakdown the whole supervised learning task in a sequence of interpretable subtasks, see Fig. 1 for an illustration of the different steps.
@@
Mission segmentation – The assessment process starts by defining the mission plan (subparagraph 6b in CS-25.571, [2]). This process consists in partitioning a mission into a structured sequence of flight segments, see Fig. 2. Each of these segments characterize different specific phases of the flight, so that the categorical variable segment $\in$ {taxi, climb, cruise, descent, approach}. It is also useful to classify segments in the ground phase (taxi) or in the flight phase (every other segment except taxi) [^1] While in rigor there exist ground segments other than taxi (e.g. take-off, braking, rotation, etc), these are not explicitly taken into account in the models as the wing PSEs don’t show any significant damage induced by these type of segments.. Accordingly, a mission flight plan is simply represented as a sequence of segments. At the same time, each segment is characterized by a vector of 12 flight parameters ${\bf P}\in\mathbb{R}^{12}$, that includes variables such as the degree of extension of the flaps in that segment, the true airspeed during the segment, the altitude at which the segment takes place, and so on (see Table 8 for details). Accordingly, within each mission, any given flight is described in terms of a sequence of segment feature vectors $\{{\bf P}_{i}\}$ with class metadata (segment class and ground/air typification).
@@
From aerodynamic loads to PSE stresses – Once the mission has been segmented and parameterized, Computational Fluid Dynamics (CFD) simulations are performed in each mission segment to obtain the corresponding aerodynamic loads. These loads are then propagated onto the structural level using Finite Element Method (FEM) simulations, so as to obtain the stresses at the different principal structural elements (PSEs) of interest (appendix 2 in CS-25.571, [2]). The resulting values will be used for the stress spectrum generation. For each PSE, we have a particular four-dimensional stress vector ${\bf S}\in\mathbb{R}^{4}$, whose elements are ${\bf S}=(1g,\Delta\text{vman},\Delta\text{vgust},\Delta{\text{turn}})$, where $1g$ is the mean equilibrium stress at the studied PSE in steady flight (i.e. the structural response under nominal flight conditions), whereas the other three variables are incremental stresses at the studied PSE caused by a standard load factor of $1.5g$ due to vertical maneuvers, vertical gusts and turns. These latter variables capture the effects of atmospheric turbulence and pilot-induced maneuvers (see Table 9 for details). The incremental stresses values depend on the modeling. To account for gusts, ESDU data or PSD models (CS-25.341 in [2]) can be used and for maneuvers military regulations specify the requirements [20]. This step is computationally intensive and must be repeated for each mission segment individually, making it one of the most resource-consuming parts of the process.
@@
From stresses to GAG and G&M cycles – With the stress data obtained in the previous step and historical flight records from a similar aircraft, two types of load cycles are derived for each mission: the so-called Ground-Air-Ground (GAG) cycle and the cycles induced by gusts and maneuvers. The GAG cycle represents the dominant stress fluctuation experienced as the aircraft transitions from ground to flight segments. This cycle typically involves the largest stress amplitude of the flight (minimum to maximum stress) and occurs once per flight. In contrast, gusts and maneuvers generate multiple smaller cycles within each flight segment. These are constructed using statistical load distributions and serve as an approximation of the fatigue loads that the aircraft will encounter across all the mission flights. Each resulting cycle is represented as a triplet: the maximum stress, the minimum stress, and the number of times that specific cycle occurs. Together, the cycles induced by gusts and maneuvers, along with the GAG cycle (derived from them) constitute the two main sources of fatigue damage—typically with the GAG cycle being the dominant contributor.
@@
Final assessment of fatigue life – Once the load cycles have been defined, the fatigue damage and corresponding life of the structural component can be estimated using material properties and local stress information. A key input in this step is the so-called S–N curve, which describes the relationship between the stress (typically the stress amplitude) and the number of cycles to failure for a given material and stress concentration factor, $k_{t}$. The factor $k_{t}$ quantifies how much the local stress is amplified due to defects or geometric features such as holes, fillets, or notches. Based on these parameters, along with the load history, the mission definition and the principal structural element (PSE), the fatigue life, quantified as number-of-flights-to-failure $N$, can be formally expressed as:
@@
A linear damage accumulation model (Miner’s rule [21]) is finally adopted to estimate fatigue damage. The core idea is that the damage caused by each stress cycle is independent and additive, and thus the total damage is simply the sum of these contributions. Under this model, the damage $\mathscr{D}$ accumulated after $n$ flights per mission is simply given by:
$\mathscr{D}=\frac{n}{N},$ (2)
such that by definition, failure occurs when the number of flights reaches the critical value $n^{*}=N$, i.e. at $\mathscr{D}=1$. Although this model simplifies the effects of plasticity and the interactions between loads, it remains widely used in aerospace applications and serves as the reference for ground truth values in this study.
@@ 2.2 Fatigue life prediction using Machine Learning
The rationale starts from assuming that $\mathscr{F}()$ in Eq. 1 is a deterministic function, and aim to approximate it using a suite of ML models. Now, after preliminary inspection it was clear that modeling $\mathscr{F}$ with a single neural network led to stagnant loss curves and suboptimal learning. This, together with the fact that including every step of the process into a single neural network hinders interpretability, led us to reformulate the solution as a sequential pipeline composed of three supervised learning (regression) tasks, such that the global task (approximate $\mathscr{F}$ in Eq. 1) can be seen as a function composition, i.e. conceptually we have $\mathscr{F}=f_{3}\circ f_{2}\circ f_{1}$, where $f_{1}$, $f_{2}$ and $f_{3}$ are nonlinear functions mapping different stages of the fatigue damage assessment pipeline described in Fig. 1. As we shall see, each of these functions are approximated by different ML and statistical models.
@@
    Formally, this task involves learning a function $f_{1}$ such that
@@
    Phase II – Damage prediction: The second phase of the pipeline estimates, from the stress data, the fatigue damage accumulated at each PSE over $n$ flights of a given mission. Since the particularities of each PSE are already encoded in the different stresses, PSE is not anymore an input variable at the stage. Now, there are two distinct damage components: $\mathscr{D}_{\text{GAG}}$, the damage caused by the Ground–Air–Ground (GAG) cycle after $n$ flights, and $\mathscr{D}_{\text{G\&M}}$, the damage induced by gusts and maneuver loads during flight segments after $n$ flights. These are the two primary contributors to fatigue degradation in metallic aircraft structures. Since they are distinct, we separate the analysis and consider two separate tasks:
    (i) The first task is that of predicting the per-flight damage caused by the Ground–Air–Ground (GAG) cycle $\overline{\mathscr{D}_{\text{GAG}}}$. This prediction uses as input variables the average stress values during ground and flight segments $\bar{\mathbf{S}}_{\text{ground}}\in\mathbb{R}$ and $\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4}$ (computed from the Phase I stress predictions), along the time the aircraft spends in flight and on the ground during a single flight, $t_{\text{ground}}$, $t_{\text{flight}}\in\mathbb{R}$ and the stress concentration factor, $k_{t}\in\mathbb{R}$ (see B for details). In principle this is a per-flight quantity and thus independent of $n$. However, although $n$ is large ($n\in[800,13{,}000]$) and the average damage is expected to asymptotically converge to the true per-flight value-as per the law of large numbers-we also add $n$ as an input variable, to account to possible deviations from the true mean due to finite samples effects. Mathematically
@@
    The total GAG-related damage after $n$ flights $\mathscr{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$.
    (ii) Likewise, the second task is that of predicting the the per-flight damage caused by the Gusts and Maneuvers (G&M) cycles $\overline{\mathscr{D}_{\text{G\&M}}}$. This prediction uses the same inputs as before, except for the ground-related variables. Mathematically, we have
@@
    The total G&M-related damage after $n$ flights $\mathscr{D}_{\text{G\&M}}=\overline{\mathscr{D}_{\text{G\&M}}}\cdot n$
    We train two different MLPs to approximate $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$. Note that the use of averaged stress values instead of segment-wise cycle estimation is motivated by the high variability of load spectra across flights and the lack of deterministic predictability of individual stress cycles. Since missions typically involve a large number of flights ($n\in[800,13{,}000]$), this formulation is justified by the Central Limit Theorem and the assumption of independent (Miner’s rule), identically distributed flights. Both models were trained using log-transformed damage values to address the wide dynamic range in the output space, and we use a Mean Absolute Error (MAE) loss function to penalize poor performance on high-damage cases, which are the most critical in fatigue life estimation. Full details on the data preparation, feature engineering, model architecture and split strategy are provided in B.
@@
    Phase III - Life prediction: The final phase computes the fatigue life $N$ defined as the number-of-flights-to-failure for a given mission and PSE. The model takes as inputs the number of flights, $n$, and the two sources of accumulated fatigue damage $\mathscr{D}_{\text{GAG}},\mathscr{D}_{G\&M}$ predicted in Phase II, such that $N=f_{3}(\mathscr{D}_{\text{GAG}},\mathscr{D}_{G\&M})$. Rather than approximating $f_{3}$ via an additional ML model, we instead apply the simpler Miner’s rule directly, as described in subsection 2.1. This linear damage accumulation model assumes failure occurs when total damage reaches unity. Accordingly, fatigue life is computed as:
    $N=f_{3}(n,\mathscr{D}_{\text{GAG}},\mathscr{D}_{G\&M})=\frac{n}{\mathscr{D}_{\text{GAG}}+\mathscr{D}_{G\&M}}.$ (3)

### v2

@@ Abstract
CI95 95% Confidence Interval
CLT Central Limit Theorem
CV Coefficient of Variation
EASA European Aviation Safety Agency
ESDU Engineering Sciences Data Unit
@@
IQR Interquartile Range
KS Kolmogorov–Smirnov
@@ 1 Introduction
Damage estimation caused by fatigue [grover1966fatigue] is a critical aspect of aeronautical engineering, as failure to predict it accurately can lead to catastrophic structural breakdowns and compromise both flight safety and operational efficiency. Likewise, fatigue damage assessment is essential for implementing predictive maintenance strategies that extend the life of aircraft and reduce unplanned downtime. Standard fatigue prediction methods rely on a convoluted and resource-intensive process that combines service history of aircraft of similar structural design with aerodynamic and Finite Element Methods simulations and cumulative damage modeling to guide expert-based decision making (CS-25.571 in [easaCS25]). The complexity of the whole process makes the development of tailored maintenance routines for new aircraft operators particularly expensive and time-consuming, involving multidisciplinary teams and a range of iterative procedures.
@@
Interestingly, the advent of Machine Learning (ML) techniques and the rapid incorporation of these tools within the aeronautical engineering community (encompassing fields from aerodynamics [linse1993identification, suresh2003lift, secco2017artificial], stress prediction [lacasa2025certification] or aeroacoustics [ALGUACIL2021116285, ruettgers2020acoustic] to airflow optimization [ramos2025transfer] –see [LeClainche2023] for a recent review–) provides a realistic opportunity to streamline and optimize the procedures for fatigue-related damage prediction, as these techniques potentially offer near-instantaneous inference of fatigue life, once properly trained. Specifically in fatigue-related problems, ML has been recently used for load spectrum reconstruction [graziani2024prediction, Valdes2017], fatigue life prediction [palczynski2022application, chen2023pinn], and crack growth modeling [Zhang2020, Younis2023]. However, most approaches treat these steps separately and rely on hybrid frameworks that combine ML with traditional tools, lacking end-to-end ML-enabled fatigue life prediction pipelines.
@@
At the same time, there is an urge of incorporating fully certifiable ML-based procedures within the aeronautical industry. As a matter of fact, a key consideration for ML in safety-critical applications is the ability to certify the predictive model. Aeronautical authorities such as the European Aviation Safety Agency (EASA) have outlined the main challenges and goals of these industries for trustworthy ML adoption [CoDANN, CoDANN2], and recent research has started to propose practical methodologies for statistical validation of ML pipelines in industrial settings [lacasa2025certification], encompassing data preparation, model training and interpretability, and multiscale uncertainty quantification.
@@
This work proposes a fully integrated and certifiable ML-based pipeline capable of predicting fatigue life at various locations of an aircraft wing, based solely on flight parameters. The pipeline mostly uses a suite of different multilayer perceptrons (MLPs) to approximate different input-output relationships within the fatigue damage assessment. An MLP is an archetypal method in supervised machine learning [goodfellow2016deep] used to infer (regress or classify) complex input-output representations $x\to y$, where $x\in\mathbb{R}^{m}$ and $y$ is usually another vector for regression tasks or an element from a discrete set in classification tasks. Often visualized as an (artificial neural) network, an MLP consists in a (deep) feedforward neural network representing an overparameterized nonlinear function $\mathscr{F}(x;{\Omega})$, where $\Omega=\{w_{k}\}$ –the set of trainable parameters– identify with the weights of the links between neurons (see [goodfellow2016deep] for details). The precise architecture of the MLP (number of neurons per layer, number of layers) as well as the precise optimization scheme to train the network are defined in a case-by-case basis.
@@
The rest of this paper is structured as follows: in Section 2, we present the full methodology: we define the physical problem under study, the dataset, the set of ML techniques and software tools, and the statistical techniques used for a thorough validation. Then, in Section 3 we present the results, which we anticipate yield a successful ML-driven fatigue life prediction tool. Finally, conclusions and future directions are discussed in Section 4.
@@ 2 Methodology
In this section, we present the methodology adopted in this work. First, in subsections 2.1 and 2.2, we describe the physical problem being modeled using machine learning techniques and explain how it is decomposed into smaller, learnable subproblems. Then, in subsection 2.3, we detail the available dataset and the data-splitting strategy used for training, which plays a crucial role in the performance of the models. Finally, in subsection 2.4, we add a few comments on the ML and statistical software libraries we have used.
@@ 2.1 Breaking down the fatigue damage assessment
The goal of this study is to develop a machine learning pipeline to predict fatigue life (expressed as the number of flights to failure) at multiple wing locations –known as principal structural elements (PSEs)–, when flying a specific mission (i.e., a flight characterized by specific flight parameters and operational objectives). In the traditional life estimation method we aim to model, all PSEs are considered non-damage-tolerant, and consequently failure is defined as the number of flights until a detectable crack initiates. Here we adopt this very same definition. In what follows we unfold the process and breakdown the whole supervised learning task in a sequence of interpretable subtasks, see Fig. 1 for an illustration of the different steps.
@@
Mission segmentation – The assessment process starts by defining the mission plan (subparagraph 6b in CS-25.571, [easaCS25]). This process consists in partitioning a mission into a structured sequence of flight segments, see Fig. 2. Each of these segments characterize different specific phases of the flight, so that the categorical variable segment $\in$ {taxi, climb, cruise, descent, approach}. It is also useful to classify segments in the ground phase (taxi) or in the flight phase (every other segment except taxi) [^1] While in rigor there exist ground segments other than taxi (e.g. take-off, braking, rotation, etc), these are not explicitly taken into account in the models as the wing PSEs don’t show any significant damage induced by these type of segments.. Accordingly, a mission flight plan is simply represented as a sequence of segments. At the same time, each segment is characterized by a vector of 12 flight parameters ${\bf P}\in\mathbb{R}^{12}$, that includes variables such as the degree of extension of the flaps in that segment, the true airspeed during the segment, the altitude at which the segment takes place, and so on (see Table 8 for details). Accordingly, within each mission, any given flight is described in terms of a sequence of segment feature vectors $\{{\bf P}_{i}\}$ with class metadata (segment class and ground/air typification).
@@
From aerodynamic loads to PSE stresses – Once the mission has been segmented and parameterized, Computational Fluid Dynamics (CFD) simulations are performed in each mission segment to obtain the corresponding aerodynamic loads. These loads are then propagated onto the structural level using Finite Element Method (FEM) simulations, so as to obtain the stresses at the different principal structural elements (PSEs) of interest (appendix 2 in CS-25.571, [easaCS25]). The resulting values will be used for the stress spectrum generation. For each PSE, we have a particular four-dimensional stress vector ${\bf S}\in\mathbb{R}^{4}$, whose elements are ${\bf S}=(1g,\Delta\text{vman},\Delta\text{vgust},\Delta{\text{turn}})$, where $1g$ is the mean equilibrium stress at the studied PSE in steady flight (i.e. the structural response under nominal flight conditions), whereas the other three variables are incremental stresses at the studied PSE caused by a standard load factor of $1.5g$ due to vertical maneuvers, vertical gusts and turns. These latter variables capture the effects of atmospheric turbulence and pilot-induced maneuvers (see Table 9 for details). The incremental stresses values depend on the modeling. To account for gusts, Engineering Sciences Data Unit (ESDU) data or Power Spectral Density (PSD) models (CS-25.341 in [easaCS25]) can be used, and for maneuvers, military regulations specify the requirements [mil8861b]. This step is computationally intensive and must be repeated for each mission segment individually, making it one of the most resource-consuming parts of the process.
@@
From stresses to GAG and G&M cycles – With the stress data obtained in the previous step and historical flight records from a similar aircraft, two types of load cycles are derived for each mission: the so-called Ground-Air-Ground (GAG) cycle and the cycles induced by gusts and maneuvers (G&M). The GAG cycle represents the dominant stress fluctuation experienced as the aircraft transitions from ground to flight segments. This cycle typically involves the largest stress amplitude of the flight (minimum to maximum stress) and occurs once per flight. In contrast, gusts and maneuvers generate multiple smaller cycles within each flight segment. These are constructed using statistical load distributions and serve as an approximation of the fatigue loads that the aircraft will encounter across all the mission flights. Each resulting cycle is represented as a triplet: the maximum stress, the minimum stress, and the number of times that specific cycle occurs. Together, the cycles induced by gusts and maneuvers, along with the GAG cycle (derived from them) constitute the two main sources of fatigue damage—typically with the GAG cycle being the dominant contributor.
@@
Final assessment of fatigue life – Once the load cycles have been defined, the fatigue damage and corresponding life of the structural component can be estimated using material properties and local stress information. A key input in this step is the so-called S–N curve, which describes the relationship between the stress (typically the stress amplitude) and the number of cycles to failure for a given material and stress concentration factor, $k_{t}$. The factor $k_{t}$ quantifies how much the local stress is amplified due to defects or geometric features such as holes, fillets, or notches. Based on these parameters, along with the load history, the mission definition and the principal structural element (PSE), the fatigue life, quantified as number-of-flights-to-failure NN, can be formally expressed as:
@@
A linear damage accumulation model (Miner’s rule [bland1946discussion]) is finally adopted to estimate fatigue damage. The core idea is that the damage caused by each stress cycle is independent and additive, and thus the total damage is simply the sum of these contributions. Under this model, the damage $\mathbb{D}$ accumulated after $n$ flights per mission is simply given by:
$\mathbb{D}=\frac{n}{N},$ (2)
such that by definition, failure occurs when the number of flights reaches the critical value $n^{*}=N$, i.e. at $\mathbb{D}=1$. Although this model simplifies the effects of plasticity and the interactions between loads, it remains widely used in aerospace applications and serves as the reference for ground truth values in this study.
@@ 2.2 Fatigue life prediction using Machine Learning
The rationale starts from assuming that $\mathscr{F}(\cdot)$ in Eq. 1 is a deterministic function, and aim to approximate it using a suite of ML models. Now, after preliminary inspection it was clear that modeling $\mathscr{F}$ with a single neural network led to stagnant loss curves and suboptimal learning. This, together with the fact that including every step of the process into a single neural network hinders interpretability, led us to reformulate the solution as a sequential pipeline composed of three supervised learning (regression) tasks, such that the global task (approximate $\mathscr{F}$ in Eq. 1) can be seen as a function composition, i.e. conceptually we have $\mathscr{F}=f_{3}\circ f_{2}\circ f_{1}$, where $f_{1}$, $f_{2}$ and $f_{3}$ are nonlinear functions mapping different stages of the fatigue damage assessment pipeline described in Fig. 1. As we shall see, each of these functions are approximated by different ML and statistical models.
@@
    Formally, this task would involve learning a function $f_{1}$ such that
@@
    Phase II – Damage prediction: The second phase of the pipeline estimates, from the stress data, the fatigue damage accumulated at each PSE over $n$ flights of a given mission. Since the particularities of each PSE are already encoded in the different stresses, PSE is not anymore an input variable at the stage. Now, there are two distinct damage components: $\mathbb{D}_{\text{GAG}}$, the total (accumulated) damage caused by the Ground–Air–Ground (GAG) cycle after $n$ flights, and $\mathbb{D}_{\text{G\&M}}$, the total (accumulated) damage induced by gusts and maneuver loads during flight segments after $n$ flights. These are the two primary contributors to fatigue degradation in metallic aircraft structures. Since they are distinct, we separate the analysis and consider two separate tasks:
    (i) The first task is that of predicting the typical ‘per-flight’ damage caused by the Ground–Air–Ground (GAG) cycle.Defining $\mathscr{D}_{\text{GAG}}$ as a latent random variable that characterises the GAG damage within a single flight, our first task is predicting the average value of this latent random variable, $\overline{\mathscr{D}_{\text{GAG}}}$, such that later we will use $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$ to estimate the total GAG damage (i.e. $\mathbb{D}_{\text{GAG}}$ is a random variable defined as the sum of $n$ independent and identically distributed random variables $\mathscr{D}_{\text{GAG}}$). The prediction of $\overline{\mathscr{D}_{\text{GAG}}}$ uses as input variables the average stress values during ground and flight segments $\bar{\mathbf{S}}_{\text{ground}}\in\mathbb{R}$ and $\bar{\mathbf{S}}_{\text{flight}}\in\mathbb{R}^{4}$ (computed from the Phase I stress predictions), along the time the aircraft spends in flight and on the ground during a single flight, $t_{\text{ground}}$, $t_{\text{flight}}\in\mathbb{R}$ and the stress concentration factor, $k_{t}\in\mathbb{R}$ (see B for details). In principle this is a per-flight quantity and thus independent of $n$. However, although $n$ is large ($n\in[800,13{,}000]$) and the average damage is expected to asymptotically converge to the true per-flight value-as per the law of large numbers, see C for more details-we also add $n$ as an input variable, to account to possible deviations from the true mean due to finite samples effects. Mathematically
@@
    As previously stated, the total GAG-related damage after $n$ flights is $\mathbb{D}_{\text{GAG}}=\overline{\mathscr{D}_{\text{GAG}}}\cdot n$.
    (ii) Likewise, the second task is that of predicting the the typical, per-flight damage caused by the Gusts and Maneuvers (G&M) cycles $\overline{\mathscr{D}_{\text{G\&M}}}$. This prediction uses the same inputs as before, except for the ground-related variables. Mathematically, we have
@@
    The total G&M-related damage after $n$ flights $\mathbb{D}_{\text{G\&M}}=\overline{\mathscr{D}_{\text{G\&M}}}\cdot n$.
    We train two different MLPs to approximate $f_{2}^{\text{GAG}}$ and $f_{2}^{\text{G\&M}}$. Note that the use of averaged stress values instead of segment-wise cycle estimation is motivated by the high variability of load spectra across flights and the lack of deterministic predictability of individual stress cycles. Since missions typically involve a large number of flights ($n\in[800,13{,}000]$), this formulation is justified by the Central Limit Theorem and the assumption of independent (Miner’s rule), identically distributed flights (see C for more details). Both models were trained using log-transformed damage values to address the wide dynamic range in the output space, and we use a Mean Absolute Error (MAE) loss function to penalize poor performance on high-damage cases, which are the most critical in fatigue life estimation. Full details on the data preparation, feature engineering, model architecture and split strategy are provided in B.
@@
    Phase III - Life prediction: The final phase computes the fatigue life NN defined as the number-of-flights-to-failure for a given mission and PSE. The model takes as inputs the number of flights, $n$, and the two sources of accumulated fatigue damage $\mathbb{D}_{\text{GAG},\mathbb{D}_{\text{G\&M}}}$ predicted in Phase II, such that $N=f_{3}(\mathbb{D}_{\text{GAG}},\mathbb{D}_{\text{G\&M}})$. Rather than approximating $f_{3}$ via an additional ML model, we instead apply the simpler Miner’s rule directly, as described in subsection 2.1. This linear damage accumulation model assumes failure occurs when total damage reaches unity. Accordingly, the total damage $\mathbb{D}=\mathbb{D}_{\text{GAG}}+\mathbb{D}_{\text{G\&M}}$ and thus fatigue life is computed as:
    $N=f_{3}(n,\mathbb{D}_{\text{GAG},\mathbb{D}_{\text{G\&M}}})=\frac{n}{\mathbb{D}_{\text{GAG}}+\mathbb{D}_{\text{G\&M}}}.$ (3)